OpenStudy (anonymous):

(f of g)(-4) f(x)=x^2+4 g(x)=2x

OpenStudy (anonymous):

F(g(x))=(2x)^2 +4 =4x^2 -4

OpenStudy (anonymous):

Sorry 4x^2 +4

OpenStudy (anonymous):

fg(x)= 4x^2 +4 fg(-4)= 64+4 =68

OpenStudy (anonymous):

That's correct

OpenStudy (anonymous):

68?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thank u soooo much

OpenStudy (anonymous):

No problem.